OpenStudy (anonymous):

The variables x and y vary inversely. Which equation relates x and y when x = -1 and y = -9?

OpenStudy (accessdenied):

An inverse relation means one variable is the other's reciprocal, at least up to some constant coefficient. \( \displaystyle xy = k \) <=> \( \displaystyle x = \frac{k}{y} \) There are three unknowns here, but we are only really interested in that constant k. If we want to find k, we just have to plug in the information we know already and evaluate.

OpenStudy (anonymous):

which on of theses makes sense A. x = y 9 B. y = -9x C. x = 9y D. y = 9

OpenStudy (accessdenied):

Sorry, I can't read those answers. Could you write them in the "Draw" window?

OpenStudy (accessdenied):

However, once you know the value of k, it should become obvious by changing k for whatever your constant is. You are given enough information to determine k with the (x = -1, y = -9) just by plugging in the values and evaluating. If you plug them into the equation, you'll have only one variable, which is easily solved for.